(A+B)2

(A+B)2

Grade:11

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below.
we can find the answer to this question simply by expanding the given equation
(A+B)2=(A+B)(A+B)
=A2+AB+BA+B2
=A2+2AB+B2